OpenStudy (anonymous):

I can't seem to take the integral from 0 to 1 of x^3 sqrt(1-x^2)dx past sin^(theta) cos^2(theta) d(theta)

OpenStudy (anonymous):

\[\int\limits_{0}^{1} x^3 \sqrt{1-x^2}dx\]

OpenStudy (anonymous):

I used x = sin (theta) dx= cos (theta) d (theta)

OpenStudy (anonymous):

I have been working on calc for 9 hours straight and my brain is mush

OpenStudy (anonymous):

try substituting u=1-x^2 du=-2x dx

OpenStudy (anonymous):

do u-sub. It's actually faster in this case

OpenStudy (anonymous):

thanks Ill try it we were working on trig substitutions so I was just stuck in that mindset

OpenStudy (anonymous):

you could do trig. But since the answer will come out the same, so why not pick the shortest way. You have plenty of examples to try trig sub